# Predictive Models in Sequential Recommendations: Bridging Performance Laws with Data Quality Insights

### v1

@@ 3.2.1 Scaling Law on Large Sequential Models
$P(N,D)\propto-[L(N,D)+$ (1)
${\alpha_{N}}ln(N)+{\alpha_{D}}ln(D)]$
$P(N,D)\propto-L(N,D)$ (2)
$=-\left[\frac{N_{c}}{N}^{\frac{\alpha_{N}}{\alpha_{D}}}+\frac{D_{C}}{D}\right]^{\alpha_{D}},$
@@ 3.3 Theorem of Performance Law
There exist $w_{3},w_{3}^{\prime},w_{4},w_{4}^{\prime}\geq 0$ such that
$d_{emb}^{w3}d^{w4}\leq Z_{t}\leq d_{emb}^{w3^{\prime}}d^{w4^{\prime}}$ (26)
@@
First, we’ll prove the left half of the inequality. This is because:
$d^{1}d_{emb}^{0}=d*0\leq\frac{d\cdot V_{n}(\sqrt{\frac{n}{2\pi e}})}{\exp(\sqrt{\frac{n}{2\pi e}})}\leq Z_{t}$ (27)
@@
$Z_{t}\leq d\cdot V_{n}(\sqrt{\frac{n}{2\pi e}})=\frac{\pi^{\frac{n}{2}}\left(\sqrt{\frac{n}{2\pi e}}\right)^{n}}{\left(\Gamma(\frac{n}{2}+1\right))}\sim k\frac{1}{\sqrt{\frac{T_{max}d_{emb}}{2}\pi e}}$ (28)
Here, we apply Stirling’s approximation $\Gamma(z+1)\sim\sqrt{2\pi z}\left(\frac{z}{e}\right)^{z}$ as the proof of the rightmost inequality. This approach enables us to decompose the loss into the form $\frac{1}{n}+\log(n)$ with a decay term, optimizing the fit for performance. Analogously to (Wu & Tang, 2024), we factorize the product into the structure of Equation 9, which represents the ultimate form of our fitting model. ∎

### v2

@@ 3.3 Theorem of Performance Law
There exist $w_{3},w_{3}^{\prime},w_{4},w_{4}^{\prime}$ such that
$log(d_{emb}^{w3}d^{w4})+$ $\frac{1}{d_{emb}^{w3}d^{w4}}\leq log(Z_{t})+\frac{1}{Z_{t}}$ (24)
$\leq log(d_{emb}^{w3^{\prime}}d^{w4^{\prime}})+\frac{1}{d_{emb}^{w3^{\prime}}d^{w4^{\prime}}}$
@@
First, assume $n=S_{max}d_{emb}$ we’ll prove the left half of the inequality. This is because:
$log(d^{0}d_{emb}^{0})+\frac{1}{d_{emb}^{0}d^{0}}=1\leq log(Z_{t})+\frac{1}{Z_{t}}$ (25)
@@
$Z_{t}\leq d\cdot V_{n}(\sqrt{\frac{n}{2\pi e}})$ $=\frac{d\pi^{\frac{n}{2}}\left(\sqrt{\frac{n}{2\pi e}}\right)^{n}}{\left(\Gamma(\frac{n}{2}+1\right))}\sim k\frac{d}{\sqrt{\frac{T_{max}d_{emb}}{2}\pi e}}$ (26)
$log(Z_{t})+\frac{1}{Z_{t}}$ $\leq O(log(d\cdot d_{emb}^{-\frac{1}{2}})+\frac{1}{d\cdot d_{emb}^{-\frac{1}{2}}})$
Here, we apply Stirling’s approximation $\Gamma(z+1)\sim\sqrt{2\pi z}\left(\frac{z}{e}\right)^{z}$ as the proof of the rightmost inequality. This requires $Z_{t}\leq 1$, which is common in the overfitting phenomenon when the model studied in this paper increases rapidly, as evidenced by the u-shape images in the experiments. This approach enables us to decompose the loss into the form $\frac{1}{n}+\log(n)$ with a decay term, optimizing the fit for performance. Analogously to (Wu & Tang, 2024), we factorize the product into the structure of Equation 7, which represents the ultimate form of our fitting model. ∎